OpenStudy (lgbasallote):

LGBADERIVATIVE: \[\frac{d^2y}{dx^2} \frac{x^2}{1+2x}\]

OpenStudy (lgbasallote):

looks complicated..

hero (hero):

Dude, no one wants to do your homework

OpenStudy (lgbasallote):

lol im checking it

OpenStudy (lgbasallote):

\[y' = \frac{2x(1+2x) - 2x^2}{(1+2x)^2}\] right?

OpenStudy (helder_edwin):

isn't there missing an equal sign or something?

OpenStudy (lgbasallote):

equal sign?

OpenStudy (lgbasallote):

i meant what's the second derivative of that thingy above...i think i may have used wrong notations

OpenStudy (zarkon):

this \[\frac{d^2y}{dx^2}=\frac{x^2}{1+2x}\] or this \[\frac{d^2}{dx^2}\left[ \frac{x^2}{1+2x}\right]\]

OpenStudy (anonymous):

Umm this derivative isn't really hard, I'm confused why you need help?

OpenStudy (helder_edwin):

y' is fine

OpenStudy (lgbasallote):

second one

OpenStudy (lgbasallote):

because if i did the first derivative right....it looks complicated

OpenStudy (anonymous):

Not to be rude, just, you usually ask more advanced stuff than this :P

OpenStudy (across):

Dude, learn the quotient rule.

OpenStudy (lgbasallote):

@nbouscal yep im trying to enhance my derivative and integral skills...im gonna need it

OpenStudy (anonymous):

Say it with me, "bottom derivative of top minus top derivative of bottom all over bottom squared"

OpenStudy (anonymous):

@nbouscal, the derivative chorus?

OpenStudy (anonymous):

Yes, this is how we learn the calculus, forget about the mathematics behind it and just learn the chanting :)

OpenStudy (anonymous):

-snickers- Very well then. :P

OpenStudy (zarkon):

(Ho D(Hi)- Hi D(Ho))/(Ho Ho)

OpenStudy (across):

\[\frac{d}{dx}\left[\frac{f(x)}{g(x)}\right]=\frac{f'(x)g(x)-f(x)g'(x)}{[g(x)]^2}\];P

OpenStudy (anonymous):

Forget \((x)\). Be Lagrangian: \[\frac{f'}{g'}=\frac{f'g-fg'}{g^2}\]

OpenStudy (anonymous):

\[y' = \frac{2x(1+2x) - 2x^2}{(1+2x)^2}\\ y' = \frac{2x^2+2x}{4x^2+4x+1}\\ y''=\frac{(4x^2+4x+1)(4x+2)-(2x^2+2x)(8x+4)}{(4x^2+4x+1)^2} \]Now simplify.

OpenStudy (lgbasallote):

is that the same as \[\Large \frac{2(1+2x)^3 - [2x(1+2x) = 2x^2](1+2x)}{(1+2x)^4}?\] if so..then i can finally get back to my advanced topics

OpenStudy (lgbasallote):

that equal sign is a minus

OpenStudy (anonymous):

first u need to find the first y' prime and the second y'' prime

OpenStudy (anonymous):

use the quotant rule 2 times

OpenStudy (anonymous):

so many ppl views your hw or work lol

OpenStudy (lgbasallote):

i think i got a different answer,,

OpenStudy (lgbasallote):

nope this isnt homework lol..this is practice

OpenStudy (anonymous):

I don't really feel like computing it out to compare our answers, I'm sure you can simplify the expressions lol

OpenStudy (lgbasallote):

lol i'll just assume they're the same

OpenStudy (anonymous):

I was attracted by the fact lgba is attempting to make his own type of derivative. I'm awaiting a formal definition.

OpenStudy (lgbasallote):

lol formal definition huh

OpenStudy (anonymous):

I'm pretty sure they are not the same, just from glancing at yours. I think you made some missteps in deriving.

OpenStudy (anonymous):

You post questions and you're like, LGBADERIVATIVE

OpenStudy (lgbasallote):

heh why not

OpenStudy (lgbasallote):

"find the derivative" was so drab for me

OpenStudy (anonymous):

confuse the reader perhaps

OpenStudy (lgbasallote):

well regardless of whether i was right or wrong..i'll now go back to d.e. i think i learned all i can from derivatives

OpenStudy (anonymous):

Interestingly, mine simplifies to: \(\dfrac{2}{(2 x+1)^3}\). Unfortunately, yours simplifies to: \(\dfrac{(6 x (x+1)+2)}{(2 x+1)^3}\) Simplifications performed by The Wolf (tm)

OpenStudy (anonymous):

I see what you did there: The Wolf and The Panda.